OpenStudy (anonymous):

multiply and simplify 8/x^2*x^4/4x ?

OpenStudy (kropot72):

\[(8\div x ^{2})\times(x ^{4}\div 4x)=(8\times x ^{4})\div (4\times x ^{3})=2x\]

OpenStudy (anonymous):

Thanks !

OpenStudy (kropot72):

yw :)